OpenStudy (anonymous):

what is 3% increase of 550,000

OpenStudy (anonymous):

First you have to find what 3% of 550,000 is. The easiest way to do that is by setting an equation up like this:

OpenStudy (anonymous):

3%=3/100

OpenStudy (anonymous):

Let's make x stand for what 3% of 550,000 is x/550,000

OpenStudy (anonymous):

Now multiply 100(x)=100x 3(550,000)=1,650,000 1,650,000/100=16,500

OpenStudy (anonymous):

so 3% of 550,000 is 16,500 now you said you need to find the 3% increase of 550,000 simply add 550,000+16,500=566,500

OpenStudy (anonymous):

Or just multiply 550,000 by 1.03…

OpenStudy (anonymous):

Thank you, I dont my answer was wrong. I multiplied .03 and 550000 and got 16500 then I added it to 550,000 and got 566500. Thank you my brain is really warped ITS FINALS WEEK- LOL THANK YOU SO MUCH

OpenStudy (anonymous):

NOT DONT BUT THOUGHT SEE MY BRAIN IS WARPED AAAAAAHHHHHHH

OpenStudy (anonymous):

Ha ha no problem!

OpenStudy (anonymous):

:)